# Average-Case Deterministic Query Complexity of Boolean Functions with Fixed Weight

### v2

@@ Abstract
We explore the average-case deterministic query complexity of boolean functions under a uniform distribution, denoted by $\mathrm{D}_{\mathrm{ave}}(f)$, the minimum average depth of zero-error decision tree computing a boolean function $f$. This measure has found several applications across diverse fields, yet its understanding is limited. We study $\mathrm{D}_{\mathrm{ave}}(f)$ of several functions, including the penalty shoot-out function, symmetric functions, linear threshold functions and the tribes functions.
We prove $\mathrm{D}_{\mathrm{ave}}(f)\leq\max\{\log\frac{\operatorname{wt}(f)}{\log n}+O(\log\log\frac{\operatorname{wt}(f)}{\log n}),O(1)\}$ for every $n$-variable boolean function $f$, where $\operatorname{wt}(f)$ denotes the weight (the number of inputs on which $f$ outputs $1$). For any $4\log n\leq m(n)\leq 2^{n-1}$, we prove the upper bound is tight up to an additive logarithmic term for almost all $n$-variable boolean functions with weight $\operatorname{wt}(f)=m(n)$.
Using Håstad’s switching lemma or Rossman’s switching lemma [Comput. Complexity Conf. 137, 2019], one can derive $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{O(w)})$ or $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{O(\log s)})$ for CNF/DNF formulas of width $w$ or size $s$, respectively. We show that, for any $w\geq\log n+\log\log n+3$, there exists a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ such that $\mathrm{D}_{\mathrm{ave}}(f)=n(1-\frac{\log n}{\Theta(w)})$. In other words, we show the criticality upper bounds $O(w)$ and $O(\log s)$ are tight up to a multiplicative $\log n$ factor, providing evidence on the tightness of the switching lemmas.
@@ 1 Introduction
The average-case deterministic query complexity of a boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ under a uniform distribution [^1] In [AD01], the average-case query complexity under distribution $\mu$ is denoted by $\mathrm{D}^{\mu}(f)$; throughout this paper, we always assume the input distribution is uniform., denoted by $\mathrm{D}_{\mathrm{ave}}(f)$, is the minimum average depth of zero-error decision trees computing $f$. This notion serves as a natural average-case analogy of the classic deterministic query complexity and has found applications in query complexity, boolean function analysis, learning algorithms, game theory, circuit complexity and percolation theory. Besides that, $\mathrm{D}_{\mathrm{ave}}(f)$ falls outside the class of polynomially-related measures including $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$ (see the summaries in [BD02, ABK16, ABK+21] and Huang’s proof of the Sensitivity Conjecture [Hua19].); $\mathrm{D}_{\mathrm{ave}}(f)$ is a measure with limited understanding. In addition, this work is also inspired by Rossman’s average-case circuit lower bounds of detecting $k$-clique on Erdős–Rényi graphs [Ros08, Ros14]. In this paper, we aim to initiate a comprehensive study on $\mathrm{D}_{\mathrm{ave}}(f)$, exploring its implications and applications.
@@ 1.1 Background
To our knowledge, Ambainis and de Wolf were the first to introduce the concept of average-case query complexity. They showed super-polynomial gaps between average-case deterministic query complexity, average-case bounded-error randomized query complexity, and average-case quantum query complexity [AD01]. In another paper, Ambainis et al. [Amb+16] studied the worst-case quantum query complexity for fixed-weight functions, and proved that almost all functions (with a fixed number of variables and fixed weight) have the same quantum query complexity (up to a constant factor).
O’Donnell, Saks, Schramm, and Servedio established the OSSS inequality [OSS+05] and the OS inequality [OS06], revealing a connection between average-case deterministic query complexity, influences, average sensitivity, and Fourier coefficients. As a result, any balanced function with $\mathrm{D}(f)\leq d$ has a variable with influence at least $1/d$. Refer to [Lee10, JZ11] for simpler proofs of the OSSS inequality. O’Donnell and Servedio [OS06] discovered the OS inequality while studying learning algorithms; the OS inequality serves as a key ingredient to learning any monotone boolean function, under the uniform distribution, in time polynomial in $n$ and the decision tree size of $f$.
In the early days of theoretical computer science, Yao [Yao77] noticed that the average-case deterministic query complexity $\mathrm{D}_{\mathrm{ave}}^{\mu}(f)$ (with respect to any distribution $\mu$) lower bounds the zero-error randomized query complexity, that is, $\mathrm{D}_{\mathrm{ave}}(f)\leq\mathrm{R}_{0}(f)$. (Moreover, Yao’s min-max principle says the maximum $\mathrm{D}_{\mathrm{ave}}^{\mu}(f)$, ranging over all distributions $\mu$, coincides with the zero-error randomized query complexity $\mathrm{R}_{0}(f)$.) For example, the zero-error randomized decision tree complexity of the recursive ternary majority function $\mathrm{Maj}_{3}^{\otimes d}$, which is bounded below by $\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{3}^{\otimes d})$, has been extensively studied [SW86, JKS03, Leo13, Mag+16]; the notion of average-case deterministic query complexity sometimes plays a role in the lower bound proof. Now the best known bound is $\Omega(2.57143^{d})\leq\mathrm{R}_{0}(\mathrm{Maj}_{3}^{\otimes d})\leq O(2.64944^{d})$ [Mag+16]. In addition, when $\mu$ is uniform, $\mathrm{D}_{\mathrm{ave}}^{\mu}(\mathrm{Maj}_{3}^{\otimes d})=\mathrm{D}_{\mathrm{ave}}(\mathrm{Maj}_{3}^{\otimes d})=O(2.5^{d})$ (Exercise 8.38 in [ODo14]).
@@
By applying the OSSS and the OS inequalities, O’Donnell et al. [OSS+05] proved that $\mathrm{R}_{0}(f)\geq\mathrm{D}_{\mathrm{ave}}^{(p)}(f)\geq(n/\sqrt{4p(1-p)})^{2/3}$ for any nontrivial monotone graph property $f$ with critical probability $p$. Here, $\mathrm{D}_{\mathrm{ave}}^{(p)}(f)$ is defined under the $p$-biased distribution. The result of O’Donnell et al. made progress on Yao’s conjecture [Yao77]: $\mathrm{R}_{0}(f)=\Omega(n)$ for every nontrivial monotone graph property; on the other hand, when $p=1/2$, Benjamini et al. proved that the lower bound $\mathrm{D}_{\mathrm{ave}}(f)\geq\Omega(n^{2/3})$ is tight up to a multiplicative logarithmic factor [BSW05].
@@
The most surprising connection (application) arose in game theory. Peres, Schramm, Sheffield, and Wilson [PSS+07] studied the random-turn HEX game, in which players determine who plays next by tossing a coin before each round. They proved that the expected playing time (under optimal play) coincides with $\mathrm{D}_{\mathrm{ave}}(f)$, where $f$ is the $L\times L$ hexagonal cells connectivity function. Using the OS inequality and the results of Smirnov and Werner on percolation [SW01], Peres et al. proved a lower bound $L^{1.5+o(1)}$ on the expected playing time on an $L\times L$ board.
@@ 1.2 Our results
In query complexity, the worst-case measures $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$ are all polynomially related. In contrast, we construct monotone balanced functions such that the polynomial relation between any two of the average-case measures $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$, $\mathrm{C}_{\mathrm{ave}}(f)$, $\mathrm{bs}_{\mathrm{ave}}(f)$ and $\mathrm{s}_{\mathrm{ave}}(f)$ does not hold. (It was already known that $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$ and $\mathrm{Q}_{\mathrm{ave}}(f)$ are not polynomially related [AD01].)
As complexity measures, $\mathrm{D}_{\mathrm{ave}}(f)$ and $\mathrm{D}(f)$ behave differently. Altering the value of $f(x)$ for a single input $x\in\{0,1\}^{n}$ may result in an $\Omega(n)$ change of $\mathrm{D}(f)$, since $\mathrm{D}(f)=n$ as long as $\operatorname{wt}(f)$ is odd. On the other hand, we prove $\mathrm{D}_{\mathrm{ave}}(f)\leq\max\{O(\log\operatorname{wt}(f)-\log\log n),O(1)\}$ for every $n$-variable boolean function $f$ (Theorem 1.1). As an application, $|\mathrm{D}_{\mathrm{ave}}(f)-\mathrm{D}_{\mathrm{ave}}(g)|<O(\log\operatorname{dist}(f,g))$; in other words, $\mathrm{D}_{\mathrm{ave}}(f)$ is Lipschitz continuous, after normalizing the Hamming distance between boolean functions. In addition, Theorem 1.1 implies $|\mathrm{D}_{\mathrm{ave}}(f)-\mathrm{D}_{\mathrm{ave}}(g)|<O(1)$ whenever $f,g$ disagree on $O(\log n)$ points at most, i.e., the Hamming distance $\operatorname{dist}(f,g)=O(\log n)$.
Put differently, Theorem 1.1 reveals a connection between $\mathrm{D}_{\mathrm{ave}}(f)$ and the zero-order Fourier coefficient $\widehat{f}(\{\emptyset\})$ (or equivalently $\mathbb{E}[f]$). Combining our upper bound with the FKN theorem [FKN02] and the KS theorem [KS03], one can show that boolean functions with Fourier weights concentrated on the low levels have small $\mathrm{D}_{\mathrm{ave}}(f)$.
@@
Next, we prove Theorem 1.1 is tight up to an additive logarithmic term for almost all fixed-weight functions. Ambainis et al. [Amb+16] showed a similar phenomenon for quantum query complexity. Besides that, Theorem 1.2 enables us to implicitly select a boolean function with fixed weight and large $\mathrm{D}_{\mathrm{ave}}(f)$, serving an ingredient in the proof of our next result (Theorem 1.3).
@@
Let $m:\mathbb{N}\to\mathbb{N}$ be a function such that $4\log n\leq m(n)\leq 2^{n-1}$. For almost all boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$ with the same weight $\operatorname{wt}(f)=m(n)$,
@@
Note that, for boolean functions with $\operatorname{wt}(f)\geq 2^{n-1}$, we have similar bounds by replacing $\operatorname{wt}(f)$ by $2^{n}-\operatorname{wt}(f)$ in (1) and (2), since $\mathrm{D}_{\mathrm{ave}}(f)=\mathrm{D}_{\mathrm{ave}}(\neg f)$ for any $f$, where $\neg f$ denotes the negation of $f$.
In addition to fixed-weight functions, we also attempt to study $\mathrm{D}_{\mathrm{ave}}(f)$ for circuits. For circuits consisting of AND, OR, NOT gates with unbounded fan-in, we show $\mathrm{D}_{\mathrm{ave}}(f)\leq 2s$, if $f$ can be computed by a circuit of $s$ gates. Note that $\mathrm{D}_{\mathrm{ave}}(\mathrm{XOR}_{n})=n$, and that $\mathrm{XOR}_{n}$ is computable by $O(n)$ gates; so the bound is almost tight.
Rossman introduced the notion of criticality, defined as the minimum value $\lambda\geq 1$ such that $\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})\geq t]<(p\lambda)^{t}$ for every $p\in[0,1]$ and $t\in\mathbb{N}$, where $\mathcal{R}_{p}$ denotes a $p$-random restriction, providing a framework to reformulate the switching lemmas. In terms of criticality, Håstad proved every width-$w$ CNF/DNF formula is $O(w)$-critical [Has86], and Rossman proved that every size-$s$ CNF/DNF formula is $O(\log s)$-critical [Ros17, Ros19]. Our next result is an attempt to prove the tightness of the switching lemmas [Has86, Ros17, Ros19], i.e., the corresponding upper bounds on criticality cannot be improved.
By applying an $(\frac{1}{2\lambda})$-random restriction as a query algorithm, we observe that $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{2\lambda})+O(1)$ for any $\lambda$-critical function $f$. As such, criticality upper bounds immediately imply upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$. For example, Håstad’s switching lemma implies that $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-1/O(w))$ for any CNF/DNF formula of width $w$ (Corollary 5.7); Rossman’s switching lemma implies that $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-1/O(\log s))$ for any CNF/DNF formula of size $s$ (Corollary 5.8).
On the other hand, if $\mathrm{D}_{\mathrm{ave}}(f)\geq n(1-1/\Omega(\lambda))$, then boolean function $f$ is $\Omega(\lambda)$-critical. In other words, lower bounds on $\mathrm{D}_{\mathrm{ave}}$ imply lower bounds on criticality. To address the tightness of the switching lemmas, we show there exist boolean functions with $\mathrm{D}_{\mathrm{ave}}(f)$ nearly matching the upper bounds (Corollary 5.7, 5.8), which are obtained straightforwardly from Håstad’s switching lemma [Has86] and Rossman’s switching lemma [Ros17, Ros19]. Our result implies that the criticality bounds $O(w)$ and $O(\log s)$ are tight up to a multiplicative $\log n$ factor.
@@
For every integer $w\in[\log n+\log\log n+3,n]$, there exists some boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ computable by a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ such that
@@
Corollary 1.4.
For every integer $w\in[\log n+\log\log n+3,n]$, there exists an $\Omega(\frac{w}{\log n})$-critical DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$.
Besides the connection with criticality, it is meaningful to understand whether the upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$ (that follow from criticality upper bounds) are tight or not. For example, consider the bound $\mathrm{D}_{\mathrm{ave}}(f)\leq n(1-\frac{1}{O(\log s)^{d-1}})$ for depth-$d$ size-$s$ $\mathsf{AC}^{0}$ circuits (Corollary 5.9). If the bound is tight, it would imply that random restriction is essentially the optimal average-case query algorithm for generic $\mathsf{AC}^{0}$ circuits. Otherwise, there exists a better query algorithm outperforming random restrictions. Either way, the answer would be interesting.
In Appendix B, we study (compute) the average-case query complexity (as well as some related average-case measures) for several functions, including the AND/OR/XOR functions, the penalty shoot-out function, symmetric functions, linear threshold functions and the tribes functions.

### v3

@@ Abstract
We study the average-case deterministic query complexity of boolean functions under a uniform input distribution, denoted by $\mathrm{D}_{\mathrm{ave}}(f)$, the minimum average depth of zero-error decision trees that compute a boolean function $f$. This measure has found several applications across diverse fields, yet its understanding is limited.
We study boolean functions with fixed weight, where weight is defined as the number of inputs on which the output is $1$. We prove $\mathrm{D}_{\mathrm{ave}}(f)\leq\max\left\{\log\frac{\operatorname{wt}(f)}{\log n}+O(\log\log\frac{\operatorname{wt}(f)}{\log n}),O(1)\right\}$ for every $n$-variable boolean function $f$, where $\operatorname{wt}(f)$ denotes the weight. For any $4\log n\leq m(n)\leq 2^{n-1}$, we prove the upper bound is tight up to an additive logarithmic term for almost all $n$-variable boolean functions with fixed weight $\operatorname{wt}(f)=m(n)$.
Håstad’s switching lemma or Rossman’s switching lemma [Comput. Complexity Conf. 137, 2019] implies $\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{O(w)}\right)$ or $\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{O(\log s)}\right)$ for CNF/DNF formulas of width $w$ or size $s$, respectively. We show there exists a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ such that $\mathrm{D}_{\mathrm{ave}}(f)=n\left(1-\frac{\log n}{\Theta(w)}\right)$ for any $w\geq 2\log n$.
@@ 1 Introduction
The average-case deterministic query complexity of a boolean function $f$ under a uniform input distribution [^1] In [AD01], the complexity under distribution $\mu$ is denoted by $\mathrm{D}^{\mu}(f)$, and $\mu$ could be arbitrary. In this paper, we assume the input distribution $\mu$ is uniform., denoted by $\mathrm{D}_{\mathrm{ave}}(f)$, is the minimum average depth of zero-error decision trees that compute $f$. This notion serves as a natural average-case analogy of the classic deterministic query complexity $\mathrm{D}(f)$ and has found applications in query complexity, boolean function analysis, learning algorithms, game theory, and percolation theory. Besides that, $\mathrm{D}_{\mathrm{ave}}(f)$ is a measure with limited understanding, since $\mathrm{D}_{\mathrm{ave}}(f)$ falls outside the class of polynomially-related measures, which includes $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, and $\mathrm{s}(f)$ (see the summaries in [BD02, ABK16, ABK+21] and Huang’s proof of the Sensitivity Conjecture [Hua19]). This work is also inspired by Rossman’s circuit lower bounds of detecting $k$-clique on Erdős–Rényi graphs in the average case [Ros08, Ros14]. Through this paper, we hope to initiate a comprehensive study on $\mathrm{D}_{\mathrm{ave}}(f)$, exploring its implications and applications.
@@ 1.1 Background
To our knowledge, Ambainis and de Wolf were the first to introduce the concept of average-case query complexity [AD01]. They showed super-polynomial gaps between average-case deterministic query complexity, average-case bounded-error randomized query complexity, and average-case quantum query complexity.
Prior to the conceptualization by Ambainis and de Wolf, average-case query complexity had been studied implicitly since the early days of computer science. Yao [Yao77] noticed that $\mathrm{D}_{\mathrm{ave}}^{\mu}(f)$ (with respect to any distribution $\mu$) lower bounds the zero-error randomized query complexity, i.e., $\mathrm{D}_{\mathrm{ave}}^{\mu}(f)\leq\mathrm{R}_{0}(f)$. Furthermore, Yao’s minimax principle says the maximum value of $\mathrm{D}_{\mathrm{ave}}^{\mu}(f)$ over all distributions $\mu$ equals $\mathrm{R}_{0}(f)$.
@@
O’Donnell, Saks, Schramm, and Servedio established the OSSS inequality [OSS+05, Lee10, JZ11]: $\mathrm{D}_{\mathrm{ave}}^{\mu_{p}}(f)\geq\frac{\mathrm{{Var}}[f]}{\max_{i}\mathrm{{Inf}}_{i}[f]}$ for any boolean function $f$, where $\mu_{p}$ is the $p$-biased distribution and $\mathrm{{Inf}}_{i}[f]$ is the influence of coordinate $i$. By applying the inequality, O’Donnell et al. [OSS+05] showed that $\mathrm{R}_{0}(f)\geq\mathrm{D}_{\mathrm{ave}}^{\mu_{p}}(f)\geq(n/\sqrt{4p(1-p)})^{2/3}$ for any nontrivial monotone $n$-vertex graph property $f$ with critical probability $p$. This result made progress on Yao’s conjecture [Yao77], which asserts that $\mathrm{R}_{0}(f)=\Omega(n)$ for every nontrivial monotone graph property. When $p=1/2$, we have $\mathrm{R}_{0}(f)\geq\mathrm{D}_{\mathrm{ave}}(f)\geq n^{2/3}$; Benjamini et al. proved that the lower bound $\mathrm{D}_{\mathrm{ave}}(f)\geq n^{2/3}$ is almost tight [BSW05].
While studying learning algorithms, O’Donnell and Servedio [OS06] discovered the OS inequality: $(\sum_{i}\hat{f}(\{i\}))^{2}\leq\mathrm{D}_{\mathrm{ave}}(f)\leq\log\mathrm{DT}_{\mathrm{size}}(f)$ for any boolean function $f$, where $\hat{f}(\cdot)$ denotes $f$’s Fourier coefficient and $\mathrm{DT}_{\mathrm{size}}(f)$ denotes the decision tree size. The OS inequality plays a crucial role in learning monotone boolean functions (under the uniform distribution).
@@
The most surprising connection (application) arose in game theory. Peres, Schramm, Sheffield, and Wilson [PSS+07] studied the random-turn HEX game, in which two players determine who plays next by tossing a coin before each round. They proved that the expected playing time (under optimal play) coincides with $\mathrm{D}_{\mathrm{ave}}(f)$, where $f$ is the $L\times L$ hexagonal cells connectivity function. Using the OS inequality and the results of Smirnov and Werner on percolation [SW01], Peres et al. proved a lower bound $L^{1.5+o(1)}$ on the expected playing time on an $L\times L$ board.
@@ 1.2 Our results
The weight of a boolean function, defined as the number of inputs on which the output is $1$, is related to its query complexity. For instance, Ambainis et al. [Amb+16] proved that the quantum query complexity of almost all $n$-variable functions with fixed weight $m$ is $\Theta\left(\frac{\log m}{c+\log n-\log\log m}+\sqrt{n}\right)$, where $c>0$ is a constant. In contrast, the hardest function with weight $m$ has quantum query complexity $\Theta\left(\left({n\cdot\frac{\log m}{c+\log n-\log\log m}}\right)^{1/2}+\sqrt{n}\right)$. Ambainis et al. [Amb+16] also proved that almost all functions with fixed weight $m\geq 1$ have randomized query complexity $\Theta(n)$ as the hardest one.
Our first result proves that $\mathrm{D}_{\mathrm{ave}}(f)\leq\log\frac{m}{\log n}+O(\log\log\frac{m}{\log n})$ for any $n$-variable boolean function $f$ with weight $m\geq 4\log n$.
@@
We prove Theorem 1.1 by designing a recursive query algorithm that attains the query complexity given in (1). The algorithm queries an arbitrary bit until the subfunction’s weight becomes sufficiently small, or more specifically, smaller than the logarithm of its input length; once this border condition is met, we invoke another algorithm which, on average, takes $O(1)$ bits to query the subfunction.
Next, we prove Theorem 1.2, complementing our first result, which says that Theorem 1.1 is tight up to a lower order term for almost all fixed-weight functions.
@@
Let $m:\mathbb{N}\to\mathbb{N}$ be a function such that $4\log n\leq m(n)\leq 2^{n-1}$. For almost all boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$ with fixed weight $\operatorname{wt}(f)=m(n)$,
@@
Remark 1.3.
For boolean functions with $\operatorname{wt}(f)\geq 2^{n-1}$, we can obtain a similar bound by replacing $f$ with $\neg f$.
Beyond fixed-weight functions, we also examine CNFs, circuits, and formulas, studying the connection between $\mathrm{D}_{\mathrm{ave}}(f)$ and criticality.
Rossman introduced the notion of criticality, defined as the minimum value $\lambda\geq 1$ such that the following property holds: $\Pr_{\rho\sim\mathcal{R}_{p}}[\mathrm{D}(f|_{\rho})\geq t]\leq(p\lambda)^{t}$ for any $p\in[0,1]$ and $t\in\mathbb{N}$. In terms of criticality, Håstad’s switching lemma says every width-$w$ CNF is $O(w)$-critical [Has86]; Rossman’s switching lemma says every size-$s$ CNF is $O(\log s)$-critical [Ros17, Ros19]; Rossman proved depth-$d$ size-$s$ $\mathsf{AC}^{0}$ circuits are $O(\log s)^{d-1}$-critical [Ros19]; Harsha et al. proved depth-$d$ size-$s$ $\mathsf{AC}^{0}$ formulas are $O(\frac{1}{d}\log s)^{d-1}$-critical [HMS23].
For any $\lambda$-critical function $f$, applying a $(\frac{1}{2\lambda})$-random restriction and then querying the resulting subfunction via its optimal decision tree yields $\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{\lambda}\right)+O\left(\sqrt{\frac{n}{\lambda}}\right)$ (Lemma 4.1 from Section 4). Hence, criticality bounds imply average-case query complexity bounds for CNFs, formulas, and circuits.
For CNFs, circuits, or formulas, it is meaningful to understand whether the upper bounds on $\mathrm{D}_{\mathrm{ave}}(f)$ are tight or not. For example, consider a $w$-CNF $f$. By Lemma 4.1 from Section 4, we have $\mathrm{D}_{\mathrm{ave}}(f)\leq n\left(1-\frac{1}{O(w)}\right)$. If the bound were indeed tight, it would suggest that the $p$-random restriction, with $p=\frac{1}{10w}$, is essentially an optimal query algorithm for generic $w$-CNFs. Otherwise, either a better query algorithm exists, or a stronger version of the switching lemma can be established. Either way, the answer would be interesting.
Along this line, we show that there exists a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ with $\mathrm{D}_{\mathrm{ave}}(f)=n(1-\frac{\log n}{\Theta(w)})$. It indicates that even if there is a better query algorithm, the room for improvement is limited when $w$ is large.
@@
For any integer $w\in[2\log n,n]$, there exists a boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ computable by a DNF formula of width $w$ and size $\lceil 2^{w}/w\rceil$ such that
@@
Lastly, we define penalty shoot-out functions in Appendix A, which are monotone balanced functions, such that the gap between $\mathrm{D}(f)$ and $\mathrm{D}_{\mathrm{ave}}(f)$ is arbitrarily large. Moreover, unlike the worst-case measures $\mathrm{D}(f)$, $\mathrm{R}(f)$, $\mathrm{C}(f)$, $\mathrm{bs}(f)$, $\mathrm{s}(f)$, which are known to be polynomially related [BD02, ABK16, Hua19, ABK+21], no such polynomial relation holds between any two of the average-case analogues [^2] These average-case counterparts are defined in the uniform distribution. $\mathrm{D}_{\mathrm{ave}}(f)$, $\mathrm{R}_{\mathrm{ave}}(f)$, $\mathrm{C}_{\mathrm{ave}}(f)$, $\mathrm{bs}_{\mathrm{ave}}(f)$, $\mathrm{s}_{\mathrm{ave}}(f)$, even for monotone balanced functions [^3] Super-polynomial gaps can be demonstrated using the threshold function [AD01], the tribes function, and $\mathrm{Maj}\circ\mathrm{AND}$, all of which are monotone. An extra trick can make them balanced: given a monotone $f$, let $g=\mathrm{Maj}(f,f^{\dagger},z)$ for $z\in\{0,1\}$, where $f^{\dagger}$ denotes $f$’s dual [ODo14]. .